# Mercury: A Computational Efficiency Benchmark for Code Large Language Models

### v2

@@ 4 Experiments
In this section, we present a series of baseline experiments to improve computational efficiency by training on Mercury-train dataset and assessing on the Mercury-eval dataset. Our empirical study encompasses 10 open-source LLMs a broad parameter spectrum from 1.3 to 34 billion. For each LLM, we compare the performance of the original model and two optimization strategies: Supervised Fine-Tuning (SFT) and Direct Preference Optimization (DPO), for their potential to optimize LLM generating functionally correct and computationally efficient code. Finally, we analyzed the underlying factors contributing to the failure of LLMs on the Mercury-eval dataset.
@@ 4.1 Baselines
Supervised Fine-Tunning. Within the SFT [6] method, an LLM undergoes additional training on a small dataset, which aims to specialize the LLM to perform better on certain tasks correlated to the training dataset. To optimize the computational efficiency performance of Code-LLMs, the most intuitive strategy is to fine-tune the Code-LLM using optimal runtime solutions. In our experimental setup, we apply the unified prompt template for each Code-LLM to ensure a fair comparison. The “pretty_content” attribute fills the <task_content> placeholder, the “prompt” attribute fills the <code_starter> placeholder, and the <code_completion> placeholder is completed with the fastest solutions. To steer Code-LLMs towards generating the intended code completion, we prepend a one-shot example to the prompt template. Appendix Figure 9 presents the prompt template.
Direct Preference Optimization. Although SFT exemplifies a straightforward approach, it is susceptible to the pitfall of catastrophic forgetting [19]. To enable LLMs to align with human preferences while preserving their functional capabilities, existing methodologies employ reinforcement learning with human preference feedback (RLHF). However, RLHF introduces additional complexities and potential instabilities, necessitating significant computing resources and additional reward model training [45, 5, 37]. DPO [33] bypasses these challenges by explicitly mapping reward functions and the optimal objective. This connection demonstrates that maximizing rewards under specific constraints can be effectively addressed through a singular training phase based on data reflecting human preferences. Consequently, DPO provides a stable, efficient, and computation-friendly alternative, removing the necessity for separate reward modeling. The DPO dataset construction and training procedure are elaborated in Appendix Section A.4.
@@ 4.2 Functional Correctness Benchmarks
HumanEval assesses the functional correctness of synthesized code derived from docstrings. It contains 164 distinct Python tasks that cover several programming areas, such as language comprehension, algorithm development, and simple mathematics [8]. MBPP has a sanitized collection of 257 entry-level Python programming problems. Each problem in this dataset consists of three components: a task description, an associated code solution, and three automated test cases to validate the code functionality [3]. Both HumanEval and MBPP harness the metric $Pass$ to measure the Code-LLMs’ functional correctness, where a task is considered solved if it passes all test cases, and the total fraction of solved passed is reported [20].
@@ 4.3 Experimental Setups
Training Data.
We use Mercury-train for model training. As for the SFT process, we nominate the fastest solution as the supervised label, then format the training data as $\langle pretty\_content,prompt,solution\_optimal\rangle$. Regarding the DPO procedure, we select the top 5 pairs of solutions that exhibit the most significant discrepancy in runtime. The training date format is $\langle pretty\_content,prompt,solution\_fast,solution\_slow\rangle$.
@@ 4.4 Empirical Results
Functional correctness is the prerequisite for evaluating computational efficiency for the code generation task. Our primary objective is to enhance the computational efficiency without compromising the functional correctness. To this end, we first introduce the existing metric $Pass$ to gauge the functional correctness [8] and then leverage $Beyond$ to provide a holistic evaluation, encompassing both computational efficiency and functional correctness. Finally, we measure the $Gap$ between $Beyond$ and $Pass$ to indicate the baseline effectiveness. These experiments aim to scrutinize the innate capabilities of cutting-edge Code-LLMs and their potential after baseline fine-tuning. Therefore, extensive parameter optimization and prompt engineering were not pursued. To deliver a comprehensive evaluation, we have further integrated the HumanEval and MBPP benchmarks as supplementary measures for appraising functional correctness [8, 3].
Functional Correctness.
Table 2 lists functional correctness benchmarks over various models, showing that larger models tend to provide better functional correctness. Except for the smallest model “deepseek-coder-1.3b-base”, DPO consistently enhances the overall $Pass$ scores across most Code-LLMs, while SFT diminishes functional correctness on the largest two Code-LLMs. These findings suggest that smaller models may struggle to integrate new knowledge while preserving their original functionality, and SFT may induce catastrophic forgetting in larger models in the pursuit of heightened computational efficiency. Moreover, it is observable that the functional correctness metrics of each model consistently diminish as the difficulty level increases, where Mercury difficulty stratification probes the model’s upper capability better than the other two benchmarks.
Table 2: Functional correctness ($Pass$) evaluation results over LLMs. The underlined values denote the top-performed approaches among the original model and the established baselines. The bolded values denote the best performance on each benchmark.
@@
[…]
Code Efficiency.
Table 3: Computational efficiency ($Beyond$) evaluation results across three difficulty levels. The bolded value indicates the top performance for each metric, while the underlined values denote the most effective approaches among the original model and the baselines.
@@
[…]
+ DPO 43.70 41.02 12.99 32.42 (-0.99) 16.33 (6.61)
[…]
+ SFT 51.37 52.71 44.28 49.39 (+1.09) 16.55 (0.16)
[…]
+ SFT 39.75 26.89 9.55 25.41 (-2.04) 12.48 (3.21)
[…]
@@
[…]
Regarding the NL2Code task, once functional correctness has been assured, attention naturally pivots to enhancing computational efficiency. As depicted in Table 3, we investigate computational efficiency metrics across a spectrum of Code-LLMs. Experiments demonstrate that DPO yields a stable enhancement in computational efficiency from models exceeding 6.7 billion parameters. In contrast, SFT detracts most $Beyond$ scores from original models, suggesting that the plain SFT may not be a feasible strategy for enhancing computational efficiency. Further analysis compares the $Gap$ between $Beyond$ and $Pass$. Since the ideal $Beyond$ should be aligned with $Pass$, it shows how much the baseline method shrinks the gap between functional correctness and computational efficiency. Our findings indicate that DPO substantially narrows $Gap$ in models larger than 15 billion parameters. However, $Gap$ tends to widen in smaller models under the same configuration. This implies that larger models possess a greater capacity to assimilate the nuanced knowledge to make strides in computational efficiency while retaining their functional correctness.
@@ 4.5 Failure Analysis
Table 4: The distribution of failure cases across Code Generation, Code Execution, and Test Case errors. E/M/H indicates Easy/Medium/Hard levels, respectively.
@@ E M H E M H E M H E M H
Table 4 provides an error breakdown of where Code-LLMs misstep during the Mercury evaluation:
Generation Errors arise from syntactical issues. The common manifestations include improper indentation, mismatched parentheses, or unexpected truncation. Fine-tuning introduces additional knowledge for Code-LLMs to adapt the Mercury convention, emphasizing standard indentation, concise code, and minimal comments. Therefore, both SFT and DPO generally reduced these errors.
Execution Errors differ from Generation Errors because they occur after the code has been successfully loaded. These errors emerge as exceptions, which could stem from various issues, such as flawed code logic, execution timeouts, memory leakage, or sandbox interruption. We observe that SFT tends to aggravate these errors on most models, whereas DPO mitigates these errors successfully.
Test Case Errors are the most prevalent errors where the code is executed without exceptions but the output fails to align with the expectation.
@@ 5 Related Work
NL2Code Evaluation currently focuses on gauging the functional correctness of generated code. As a pioneer, CodeBLEU [34] adapts the BLEU [32] metric into code generation. However, given the abstract nature of programming languages, distinct code can express the equivalent semantics, prompting subsequent benchmarks to harness test cases instead of the similarity measurement. For example, HumanEval [8] and MBPP [3] consist of hand-written Python programming tasks and corresponding test cases. On the note of enhancing language inclusiveness, ODEX [41] integrates multiple natural languages, while MBXP [2] extends the benchmarks to cater to a variety of programming languages, promoting polyglot code generation evaluation. Recent benchmarks have also begun to consider more aspects beyond functional correctness. For instance, the benchmark DS-100 [21] dives deeply into the data analysis scenarios, and CodeGen [30] contributes a benchmark for multi-turn code generation. For security-oriented code generation, SecurityEval [36] offers a concentrating benchmark on mining the vulnerability of generated code.
@@ 6 Limitations
In this work, we measure computational efficiency under the assumption that the code runtime is uniformly distributed. The simplification allows computational efficiency evaluation via limited solution samples. However, the distribution of code runtime in real-world scenarios is more intricate, which may call for more solution samples to support more precise modeling. Additionally, the presence of data contamination during the model training phase compromises the precision of the Mercury benchmark to reflect the performance of tainted models [17]. To mitigate this issue, we will update our benchmark via our open-sourced data collection framework to import new tasks dynamically, thus laying the groundwork for more detailed investigations in subsequent studies.
@@ 7 Conclusion
In this work, we introduced Mercury, the first computational efficiency benchmark for NL2Code evaluation. Unlike prior work that focused on functional correctness, our benchmark highlights the importance of computational efficiency. By crafting dedicated test case generators and sampling ground-truth solutions across different difficulty levels from Leetcode, we have developed a comprehensive and rigorous Code-LLM evaluation frame. We evaluated leading Code-LLMs against benchmarks and found that even though these models are proficient in generating functionally correct code, there is still considerable space for computational efficiency improvement. As Code-LLMs become more widely used, computational efficiency determines factual productivity, where Mercury can gauge the vital metric. As a commitment to ongoing research and to foster further innovation in this area, we have open-sourced the Mercury dataset collection framework, laying the groundwork for future advancements in the field.
@@ Appendix A Appendix
A.1 Definition of Mercury Fields
@@ A.2 Mercury Data Distribution and Customized Data Structures
Table 6: Mercury-eval encompasses 256 tasks, the difficulty of which has been balanced for model evaluation. Mercury-train comprises the remaining 1,633 tasks for model training.
@@ A.3 Sandbox Details
Each executed code within the sandbox is subject to certain constraints to ensure fair utilization of resources and to prevent any single program from monopolizing the system. Specifically, there are two primary constraints: a time limit and a memory limit. The time limit restricts how long the code can execute before being forcibly terminated, thereby ensuring that no infinite loops or excessively long computations negatively impact the availability of the sandbox. The memory limit caps the amount of RAM that a process can consume. This measure prevents a single program from exhausting the server’s memory resources, which could lead to a denial of service for other users or processes sharing the same infrastructure.
@@
To mitigate harmful activities such as unauthorized data access or data exfiltration, the sandbox imposes strict input/output (IO) restrictions. These restrictions can include limitations on the size and type of files that can be read from or written to the disk and restrictions on the use of network sockets for sending or receiving data. By controlling the IO operations, the sandbox can prevent many common exploits and ensure that the code running within it cannot interfere with other processes or the host system.
@@
The sandbox employs an isolated file system to provide a safe execution environment for the code. This means that the process running in the sandbox has its virtual file system, which is separated from the host’s file system. The isolated nature of this file system ensures that even if a process within the sandbox attempts to modify or delete files, these changes will not affect the host system or other sandboxes. It acts as a layer of security, protecting the host from potential threats and maintaining the integrity of the overall system.
@@
In the context of the sandbox environment, single-threaded evaluation is the process of executing code using only one thread of execution. This approach simplifies resource management and timing measurements, avoiding the complexities associated with multi-threaded execution, such as synchronization, race conditions, and deadlocks. Single-threading can be particularly important in educational or testing settings where reproducibility and fairness are critical, ensuring that every student or user’s code is evaluated with the same computational resources.
@@
Figure 5: Sandbox Execution Pipeline. 1) Test Case Generation. We first employ the corresponding test case generator for each task to produce a comprehensive set of test cases for the subsequent evaluation. 2) Context Initialization. To prevent any unexpected code behavior, the sandbox environment is meticulously reinitialized for each new task. This phase ensures that all the common libraries required for executing the solution are loaded. 3) Solution Instantiation. The solution under evaluation will be encapsulated as a solution class. 4) Test Case Evaluation. Each test case the generator provides will be rigorously executed against the solution. A solution must successfully pass all the test cases to be deemed valid. 5) Clean up. The final stage involves the sandbox dutifully clearing the namespace environment and the temporary directory. Mercury records the time consumed during the stage of Solution instantiation and Test Ease Evaluation as the primary metric for assessing computational efficiency.

### v3

@@ 4 Experiments
In this section, we present a series of baseline experiments to improve code efficiency by training on Mercury-train dataset and assessing on the Mercury-eval dataset. Our empirical study encompasses 10 open-source LLMs with a broad parameter spectrum from 1.3 to 34 billion. For each LLM, we compare the performance of the original model and two optimization strategies, Supervised Fine-Tuning (SFT) and Direct Preference Optimization (DPO), for their potential to optimize LLM generating functionally correct and computationally efficient code. Finally, we analyzed the underlying factors contributing to the failure of LLMs on the Mercury-eval dataset.
@@ 4.1 Baselines
Supervised Fine-Tunning. Within the SFT [6] method, an LLM undergoes additional training on a small dataset, which aims to specialize the LLM to perform better on certain tasks correlated to the training dataset. To optimize the code efficiency performance of Code LLMs, the most intuitive strategy is to fine-tune the Code LLM using optimal runtime solutions. In our experimental setup, we apply a unified prompt template for each Code LLM to ensure a fair comparison. The “pretty_content” attribute fills the <task_content> placeholder, the “prompt” attribute fills the <code_starter> placeholder, and the <code_completion> placeholder is completed with the fastest solutions. To steer Code LLMs towards generating the intended code completion format, we prepend a one-shot example to the prompt template. Appendix Figure 9 presents the prompt template.
Direct Preference Optimization. Although SFT exemplifies a straightforward approach, it is susceptible to the pitfall of catastrophic forgetting [20]. To enable LLMs to align with human preferences while preserving their functional capabilities, existing methodologies employ reinforcement learning with human preference feedback (RLHF). However, RLHF introduces additional model complexities and potential instabilities, necessitating significant computing resources and extra reward model training [48, 5, 40]. DPO [36] bypasses these challenges by explicitly mapping reward functions and the optimal objective. This connection demonstrates that maximizing rewards under specific constraints can be effectively addressed through a singular training phase based on data reflecting human preferences. The DPO training procedure is elaborated in Appendix Section A.4.
@@ 4.2 Functional Correctness Benchmarks
HumanEval assesses the functional correctness of synthesized code derived from docstrings. It contains 164 distinct Python tasks that cover several programming areas, such as language comprehension, algorithm development, and simple mathematics [8]. MBPP has a sanitized collection of 257 entry-level Python programming problems. Each problem in this dataset consists of three components: a task description, an associated code solution, and three automated test cases to validate the code functionality [3]. Both HumanEval and MBPP harness the metric Pass to measure the Code LLMs’ functional correctness, where a task is considered solved if the given solution passes all test cases, and the total fraction of solved tasks is reported as $Pass=N_{solved}/N_{total}$ [21].
@@ 4.3 Experimental Setups
Training Data. We use Mercury-train for model training. As for the SFT process, we nominate the fastest solution as the supervised label, then format the training data as $\langle pretty\_content,prompt,solution\_optimal\rangle$. Regarding the DPO procedure, we select the top 5 pairs of solutions that exhibit the most significant discrepancy in runtime. The training date format is $\langle pretty\_content,prompt,solution\_fast,solution\_slow\rangle$.
@@ 4.4 Empirical Results
Functional correctness is the prerequisite for evaluating code efficiency for the code generation task. Our primary objective is to enhance the code efficiency without compromising the functional correctness. To this end, we first introduce the existing metric Pass to gauge the functional correctness [8] and then leverage Beyond to provide a holistic evaluation, encompassing both code efficiency and functional correctness. Finally, we measure the Gap between Beyond and Pass to mirror the baseline ability for improving efficiency while preserving correctness. These experiments aim to investigate the innate capabilities of cutting-edge Code LLMs and their potential after baseline fine-tuning. Therefore, extensive parameter optimization and prompt engineering were not pursued. To deliver a comprehensive evaluation, we have further integrated the HumanEval and MBPP benchmarks as supplementary measures for appraising functional correctness [8, 3].
Figure 3: The horizontal axis represents the score for functional correctness, while the vertical axis indicates the score for code efficiency. The left figure illustrates the performance of the baseline model, whereas the right one depicts the performance after DPO tuning. Model points located nearer to the diagonal line exhibit a more equitable balance between functionality and efficiency.
Functional Correctness. Table 2 lists Pass scores over various Code LLMs, showing that larger models tend to provide better functional correctness. Except for the smallest model “deepseek-coder-1.3b-base”, DPO invariably enhances the overall Pass scores across most Code LLMs, while SFT diminishes functional correctness on the largest two Code LLMs. These findings suggest that smaller models may struggle to integrate new knowledge while preserving their original functionality, and SFT may induce catastrophic forgetting in the pursuit of heightened code efficiency. Moreover, it is evident on Mercury that Pass scores of each model consistently decline as the difficulty level increases, indicating that the Mercury difficulty stratification is effective at probing the upper limitation of each Code LLM compared to the auxiliary benchmarks.
Table 2: Functional correctness (Pass) evaluation results. The underlined values denote the top-performed approaches among the original model and baselines. The bolded values denote the best performance on each benchmark. We sample one solution for each task to calculate pass score.
@@
+ DPO 65.9 75.2 83.9 68.4 63.2 70.6 (+18.2)
Table 3: Code efficiency (Beyond) evaluation results across three difficulty levels. The bolded value indicates the top performance for each metric, while the underlined values denote the most effective approaches among the original model and the baselines. In our experiment, we sample 5 solutions for each task to calculate Beyond score.
@@
+ DPO 46.91 42.27 16.78 35.21 (-0.41) 9.64 (-0.21)
starcoder2-3b 43.55 41.91 15.21 33.40 9.72
+ SFT 44.64 42.10 15.72 34.01 (+0.61) 14.04 (+4.31)
+ DPO 43.70 41.02 12.99 32.42 (-0.99) 16.33 (+6.61)
deepseek-coder-6.7b-base 48.80 51.16 45.11 48.29 16.40
+ SFT 51.37 52.71 44.28 49.39 (+1.09) 16.55 (+0.16)
+ DPO 56.25 52.35 40.62 49.70 (+1.41) 18.73 (+2.34)
starcoder2-7b 50.23 51.29 20.25 40.37 10.95
+ SFT 42.21 44.02 21.09 35.61 (-4.77) 15.80 (+4.84)
+ DPO 53.52 51.41 17.35 40.56 (+0.18) 17.41 (+6.46)
CodeLlama-7b-hf 42.55 30.99 8.88 27.45 9.27
+ SFT 39.75 26.89 9.55 25.41 (-2.04) 12.48 (+3.21)
+ DPO 54.14 34.48 11.10 33.29 (+5.84) 10.46 (1.19)
CodeQwen1.5-7B 51.11 53.56 39.03 47.78 15.35
@@
+ DPO 78.55 60.95 51.94 63.94 (+21.54) 8.01 (-7.47)
Code Efficiency. Regarding the NL2Code task, once functional correctness has been assured, attention naturally pivots to enhancing code efficiency. As depicted in Table 3, we investigate code efficiency metrics across a spectrum of Code LLMs. Experiments demonstrate that DPO yields a stable enhancement in code efficiency from models exceeding 6.7B parameters. In contrast, SFT detracts most Beyond scores from original models, suggesting that the plain SFT may not be a feasible strategy for enhancing code efficiency. Further analysis compares the Gap between Beyond and $Pass$. Since the ideal Beyond should be aligned with Pass (where the LLM-generated solution is correct and faster than all historical solutions), it shows how much the baseline method shrinks the gap between functional correctness and code efficiency. Our findings indicate that DPO substantially narrows Gap in models larger than 15B parameters. However, Gap tends to widen in smaller models under the same configuration. This implies that larger models possess a greater capacity to assimilate the nuanced knowledge to make strides in code efficiency while retaining their functional correctness.
@@ 4.5 Failure Analysis
Table 4 provides an error breakdown of where Code LLMs misstep during the Mercury evaluation:
(1) Generation Errors arise from syntactical issues. The common manifestations include improper indentation, mismatched parentheses, or unexpected truncation. Fine-tuning introduces additional knowledge for Code LLMs to adapt the Mercury convention, emphasizing standard indentation, concise code, and minimal comments. Therefore, both SFT and DPO generally reduced these errors.
(2) Execution Errors differ from Generation Errors because they occur after the code has been successfully loaded. These errors emerge as exceptions, which could stem from various issues, such as flawed code logic, execution timeouts, memory leakage, or sandbox interruption. We observe that SFT tends to aggravate these errors on most models, whereas DPO mitigates these errors successfully.
(3) Test Case Errors are the most prevalent errors where the code is executed without exceptions, but the output fails to align with the expectation.
Table 4: The distribution of failure cases across Code Generation, Code Execution, and Test Case errors. E/M/H indicates Easy/Medium/Hard levels, respectively. We sample 5 solutions for each task, so there are $256*5=1280$ solutions in total for each model.
@@ 5 Related Work
NL2Code Evaluation currently focuses on gauging the functional correctness of generated code. As a pioneer, CodeBLEU [37] adapts the BLEU [35] metric into code generation. However, given the abstract nature of programming languages, distinct code can express the equivalent semantics, prompting subsequent benchmarks to harness test case fuzzing instead of the similarity measurement. For example, HumanEval [8] and MBPP [3] consist of hand-written Python programming tasks and corresponding test cases. On the note of enhancing language inclusiveness, ODEX [43] integrates multiple natural languages, while MBXP [2] extends the benchmarks to cater to a variety of programming languages, promoting polyglot code generation evaluation. Recent benchmarks have also begun to consider more aspects beyond functional correctness. For instance, the benchmark DS-100 [22] dives deeply into the data analysis scenarios, and CodeGen [32] contributes a benchmark for multi-turn code generation. For security-oriented code generation, SecurityEval [39] offers a concentrating benchmark on mining the vulnerability of generated code.
@@ 6 Limitations
In this work, we measure code efficiency under the assumption that the code runtime is uniformly distributed. The simplification streamlines code efficiency evaluation via limited solution samples. However, the distribution of code runtime in real-world scenarios is more intricate, which may call for more solution samples to support more precise modeling. Additionally, the presence of data contamination during the model training phase compromises the precision of the Mercury benchmark to reflect the performance of tainted models [18]. To mitigate this issue, we will update our benchmark via our open-sourced data collection framework to import new tasks dynamically, thus laying the groundwork for more detailed investigations in subsequent studies.
@@ 7 Conclusion
In this work, we introduced Mercury, the first code efficiency benchmark for NL2Code evaluation. Unlike prior work that focused on functional correctness, our benchmark highlights the importance of code efficiency. By crafting dedicated test case generators and sampling ground-truth solutions across all difficulty levels from Leetcode, we have developed a comprehensive and rigorous Code LLM evaluation frame. We evaluated leading Code LLMs against benchmarks and found that even though these models are proficient in generating functionally correct code, there is still considerable space for code efficiency improvement. As Code LLMs become more widely used, code efficiency determines factual productivity, where Mercury can gauge the vital metric. As a commitment to ongoing research and to foster further innovation in this area, we have open-sourced the Mercury dataset collection framework, laying the groundwork for future advancements in the field.
Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes] The abstract and introduction include the main contributions and the research scope.
    2. (b)
      Did you describe the limitations of your work? [Yes] See Section 6.
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [N/A] Given that our work is largely technical and does not engage with societal systems directly, it is unlikely to have negative societal repercussions.
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes] The paper conforms to the ethics review guidelines.
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [N/A] The paper does not include any theoretical results.
    2. (b)
      Did you include complete proofs of all theoretical results? [N/A] The paper does not include any theoretical results.
  3. 3.
    If you ran experiments (e.g. for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] We release our dataset on HuggingFace and our code on GitHub. See the abstract footnote.
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] See Section 4.3.
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [Yes] See Section 4.
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] See Section 4.3 and Appendix Section A.11.
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [Yes] See References.
    2. (b)
      Did you mention the license of the assets? [Yes] See Section A.13
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [Yes] See the abstract footnote.
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [Yes] See Section A.13
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [Yes] See Section A.13
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [N/A]
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A]
@@ Appendix A Appendix
A.1 Dataset Nutrition Labels
@@ A.2 Mercury Data Distribution and Customized Data Structures
Table 6: Mercury-eval encompasses 256 tasks, the difficulty level of which has been balanced for model evaluation. Mercury-train comprises the remaining 1,633 tasks for model training.
@@ A.3 Sandbox Details
Each executed code within the sandbox is subject to certain constraints to ensure fair utilization of resources and to prevent any single code from monopolizing the system resource. Specifically, there are two primary constraints: a time limit and a memory limit. The time limit restricts how long the code can execute before being forcibly terminated, thereby ensuring that no infinite loops or excessively long computations negatively impact the availability of the sandbox. The memory limit caps the amount of RAM that a process can consume. This measure precludes a single code from exhausting the memory resources, which could lead to a denial of service for subsequent codes. In our experiment settings, the timeout limit is 30 seconds, and the memory limit is 2048 MB for each solution execution.
@@
To mitigate harmful activities such as unauthorized command execution or data exfiltration, the sandbox imposes strict Input/Output (IO) restrictions. These restrictions include limitations on reading from or writing to the disk and restrictions on the use of network sockets for sending or receiving data. By controlling the IO operations, the sandbox can prevent many common vulnerabilities and ensure that the code runs without interfering with other processes of the host system.
@@
The sandbox employs an isolated file system to provide a safe execution environment for the code. This means that the process running in the sandbox has its virtual file system, which is separated from the host’s file system. The isolated nature of this file system ensures that even if a process within the sandbox attempts to modify or delete files, these changes will not affect the host system or other sandboxes. It acts as a security layer, protecting the host from potential threats and maintaining the integrity of the overall system.
@@
Single-threaded evaluation refers to executing code using a sole thread of execution, thereby simplifying resource management and timing assessments, and mitigating the intricacies linked with multi-threaded execution, such as synchronization issues, race conditions, and potential deadlocks. This mode of operation is especially important in testing environments where reproducibility and fairness are paramount, ensuring that each piece of code is evaluated using identical computational resources.
@@
Figure 5: Sandbox Execution Pipeline. 1) Test Case Generation. We first employ the corresponding test case generator for each task to produce a comprehensive set of test cases for the subsequent evaluation. 2) Context Initialization. To prevent any unexpected code behavior, the sandbox environment is meticulously reinitialized for each new task. This phase ensures that all the common libraries required for executing the solution are loaded. 3) Solution Instantiation. The solution under evaluation will be encapsulated as a solution class. 4) Test Case Evaluation. Each test case the generator provides will be rigorously executed against the solution. A solution must successfully pass all the test cases to be deemed valid. 5) Clean up. The final stage involves the sandbox dutifully clearing the namespace environment and the temporary directory. Mercury records the time consumed during the stage of Solution instantiation and Test Ease Evaluation as the primary metric for assessing code efficiency.